OpenStudy (anonymous):

Calculate the number of moles of Cl atoms in 1.81×10^24 formula units of magnesium chloride,MgCl2.

OpenStudy (anonymous):

in 1 mole there are = 6.02*10^23 particles(atoms, molecules, ions, formula units..) one formula unit of MgCl2 contains 2 Cl atoms(ions in this case).. so the, 1.81*10^24 MgCl2 should containt 2*(1.81*10^24) atoms...2*1.8*10^24 = 3.6 * 10^24. if 1 mole contains 6.02*10^23 particles.. then x mole contains 3.6*10^24 particles(in our case it Cl atoms or ions) by solving this we find approximately 6 moles of Cl(atoms or ions..)